OpenStudy (anonymous):

Find the perimeter of the square. length = 15x - 42 width = 8x

OpenStudy (jiteshmeghwal9):

Since length=15x-42 width=8x Perimeter=2(L+B) =2((15x-42)-(8x))

OpenStudy (anonymous):

P = 2(l) * 2 (w)

OpenStudy (jiteshmeghwal9):

& ur question is wrong if u have an square than why u have diffrent length & width ???

OpenStudy (jiteshmeghwal9):

a square *

OpenStudy (anonymous):

me soo stupid. sory. plz continue.

OpenStudy (jiteshmeghwal9):

u should solve 2((15x-42)+(8x))

OpenStudy (jiteshmeghwal9):

@rmaker408 :)

OpenStudy (anonymous):

i got this: 46x - 84

OpenStudy (anonymous):

IF ITS SQUARE , THEN LENGTH =WIDTH SO 15x-42 = 8x find x from here. then side = 8x and perimeter = 4*(8x)

OpenStudy (anonymous):

@jiteshmeghwal9 i don't understand how to solve this. can you help me plz??

OpenStudy (jiteshmeghwal9):

2(15x-42+8x) 2(23x-42) 2(23x)-2(42)

OpenStudy (anonymous):

which is 46x + 84. is that the answer?

OpenStudy (anonymous):

the question isn't wrong, you'll get a numerical answer, see my comment, jitesh.

OpenStudy (jiteshmeghwal9):

Ok ! cn u prove me 5x-42=8x

OpenStudy (anonymous):

15x and yes. the answer is -6.

OpenStudy (anonymous):

@jiteshmeghwal9 and @gohangoku58 you two are geniuses. or is the plural of genius, genii?

OpenStudy (anonymous):

for square all sides are equal 15x - 42 =8x

OpenStudy (jiteshmeghwal9):

Okay!

OpenStudy (anonymous):

Thanks to everyone who helped me. I really appreciate it. :D

OpenStudy (anonymous):

so did u get the answer ? perimeter = ?